OpenStudy (anonymous):

Let p and q be real numbers such that p not equal0 , p^3 not equal q and p^3 not equal -q . if a and b are the non zero complex numbers satifying a + b = -p and a^3 + b^3 = q then a quadratic equation having roots a/b and b/a as its roots is?

OpenStudy (anonymous):

With these roots equation form will be: \[(bx-a)(ax-b) = 0\] \[ab x^2 - (a^2 + b^2)x + ab = 0\]

OpenStudy (anonymous):

a and b are here complex numbers..

OpenStudy (anonymous):

The answer shiuld be (p^3+ q)x^2 - (p^3 - 2q) x + (p^3 + q) =0

OpenStudy (anonymous):

So we have to prove: \[ab = p^3 + q\] \[a^2 + b^2 = p^3 - 2q\]

OpenStudy (anonymous):

\[a^3+b^3=(a+b)((a+b)^2-3ab)\]

OpenStudy (anonymous):

\[ab = \frac{p^3 + q}{3p}\]

OpenStudy (anonymous):

\[a^2 +b^2 = (a+b)^2 - 2ab\]

OpenStudy (anonymous):

\[a^2+b^2=\frac{p^3-2q}{3p}\]

OpenStudy (anonymous):

Yep..

OpenStudy (anonymous):

we are done

OpenStudy (anonymous):

Yes we are done.. There was denominator too.. I could not guess as it became 0 after solving..

OpenStudy (anonymous):

Not 0 but 1..

OpenStudy (anonymous):

\(p\neq0\) we can multiply by \(3p\)

OpenStudy (anonymous):

Yes.. That is what I meant..

OpenStudy (anonymous):

It is given in question...

OpenStudy (anonymous):

@Yahoo! you have the values can you now substitute and solve ??

OpenStudy (anonymous):

@mukushla @shubhamsrg @waterineyes .. ab = p^2 - q /3

OpenStudy (ajprincess):

 Is it k nw?